32. Do you have a great deal of confidence in the advice given to Suppose a random sample of n-37 adults in a health insurance s great deal of confidence in their M.D program are asked about their confidence in the medical advice thelir doctors give (a) Is the normal approximation to the proportion p rin valid(Round your answers to two decimal places) It n-q- possible to use the normal approximation to the binomial distribution because n p , which isDs (b) Find the values of pp and op (Round your answer for op to three decimal places.) Pp
Answer #1

Given,

p = 35% = 0.35

n = 37

a)

It is possible to use the normal approximation to Binomial distribution because n*p =37*0.35 12.95 , which is > 5 and n*q =37*(1-0.35) = 24.05 , which is > 5

b)

mu _{hat{p}} = p = 0.35

ơp = VIp(1-p) = V10.35 * (1-0.35) = 0.078 /p(1-p) _ . |0.35 * (1-0.35) Ơnー